# Optimizations of electronic stopping cross section measurements using the backscattering method

### v1

@@ Abstract
The accurate determination of electronic stopping cross sections is essential for advancing nuclear science and materials research. In this work, we present an optimized methodology for stopping power measurements using the backscattering technique, with a focus on minimizing both random and systematic uncertainties. By systematically analyzing the sources of uncertainty and refining the experimental geometry, we establish a robust framework for high-precision measurements. Our approach was applied to the stopping power of helium ions in gold thin films, demonstrating excellent agreement with reference models such as SRIM and ICRU-49. The results reveal that carefully chosen measurement angles can effectively balance statistical precision and systematic accuracy, achieving total uncertainties below 3% across a wide energy range. This study provides a refined strategy for stopping power determination, offering valuable improvements for ion beam analysis and benchmarking theoretical models.
@@
Stopping cross section , Backscattering method , Uncertainty optimization , Systematic uncertainty
@@ 1 Introduction
The electronic stopping power of ions in matter plays a crucial role in nuclear science and its applications. Although theoretical understanding of the physical processes involved in energy loss phenomena is advancing [1, 2, 3], experimental works focusing on traceable experimental measurements become crucial to benchmark models. Accurate experimental data also help to improve semiempirical approaches to produce tables that are used in simulations targeting a wide range of applications.
In this work, we demonstrate the optimization of the electronic stopping power measurements to minimize the random and systematic uncertainties in the backscattering method. Our approach consists of estimating the sources of uncertainties and improving the method by choosing the best geometry that favors the reduction of uncertainties.
@@ 2.1 The back-scattering method
There are several approaches to measuring stopping power using backscattering data [4]. In this work, we adopt the method based on determining the peak width in thin-film measurements, as described in [5]. Fig. 1 shows a typical spectrum of a pure metallic film deposited on a light substrate. The peak is formed by all scattered ions from the surface down to the film-substrate interface. Its width ($\Delta E$) is the difference between the energy of ions scattered by the film atoms at the surface ($K\cdot E_{0}$, with $K$ being the kinematic factor and $E_{0}$ the incidence beam energy) and the exit energy of ions re-emerging after being scattered by film atoms at the interface ($E_{out}$). Assuming a homogeneous stopping in the test film along both, the entry and exit paths, the energy difference can be written as:
@@
The peak width can be determined by adding the energy loss during the path of the ion into the film and the energy loss during its exit from a thin film of thickness $t$. Assuming homogeneous stopping power along both the entry and exit paths, this expression can be written as:
@@
being the stopping cross section in the $j$ path, either the entry (index $in$) or the exit (index $out$) of the film. $\theta_{in}$ and $\theta_{out}$ represent the angles between the ion’s entry and exit paths and the surface normal, respectively. $N$ is the material atomic density.
The fundamental principle of this methodology is that by measuring the peak width in two different geometrical configurations (i.e. two pairs of values of $\theta_{in}$ and $\theta_{out}$) for the same incident energy, one can obtain a system of equations. The solutions for $[\epsilon_{in}]$ and $[\epsilon_{out}]$ yield the stopping cross section at two different mean energies. In matrix formalism, this can be expressed as:
@@
where index $A$ and $B$ denote the two different measurements. To better constrain the problem, we impose the conditions $\theta_{in}^{A}=\theta_{out}^{B}=\Theta_{1}$ and $\theta_{in}^{B}=\theta_{out}^{A}=\Theta_{2}$. This assumption is necessary to reduce the number of variables, thereby enhancing result visualization and simplifying the experimental procedure. It is important to note that the factor $K$ implicitly depends on the scattering angle, which remains fixed throughout the analysis. As a consequence, the experimental procedure involves tilting the sample in such a way that, in the first measurement, the incidence angle is $\Theta_{1}$ and the exit angle is $\Theta_{2}$, while in the second measurement, the angles are reversed, with the incidence angle set to $\Theta_{2}$ and the exit angle to $\Theta_{1}$.
There are different approximations to determine the mean energy at the entry and exit paths of the film [5]. Here, we adopt the following approximations:
@@
with $\Delta\bar{E}=(\Delta E^{A}+\Delta E^{B})/2$. This is an alternative definition based on the symmetrical mean energy approximation proposed in [5], which is more convenient for an experimental approach.
@@ 2.2 Accounting for random uncertainties
From Eq. 3, we can propagate the uncertainties in the determination of $\Delta E_{j}$ to the values of $[\epsilon_{j}]$ using the definition of a covariance matrix:
@@
$X=\begin{pmatrix}\frac{K\cdot t}{\cos(\Theta_{1})}&\frac{t}{\cos(\Theta_{2})}\\ \\ \frac{K\cdot t}{\cos(\Theta_{2})}&\frac{t}{\cos(\Theta_{1})}\\ \end{pmatrix}$ (6)
@@
and the outer diagonal results in the covariance between the $[\epsilon]_{in}$ and $[\epsilon]_{out}$ values, which is used to calculate the correlation factor r as:
@@ 2.3 Accounting for systematic error
To estimate the systematic uncertainties, we tested Eqs. 4 and 8 against the SRIM curve. The peak width, $\Delta E$, has been computed directly using the integral (as in Eq. 12) form and SRIM data, while the methodology described in Section 2.1 is applied to determine $[\epsilon]$ using Eq. 8. These values are then compared to the SRIM values themselves interpolated at the mean energies given by Eq. 4. The differences between the interpolated and the output of the matrix calculation are considered as systematic errors. This accounts for systematic errors inserted either by the surface approximation or by the calculation of the mean energy in the entry and exit paths.
[…]
@@
with $s({[\epsilon]_{j}})$ being the systematic error and $\textbf{SRIM}(\bar{E_{j}})$ being the SRIM interpolation at the energy $\bar{E_{j}}$. $j$ is an index that denotes $in$ or $out$.
@@ 2.4 Finding the best geometry
With the definitions of random and systematic uncertainties for the case of determination of $[\epsilon]$, the total uncertainty can be defined as:
@@
Note that the common factor in the statistical uncertainty of $[\epsilon]_{in}$ and $[\epsilon]_{out}$, as in Eq. 9 and 10, can be written as the function $f(\Theta_{1},\Theta_{2})$:
@@
which diverges when $\Theta_{1}\approx\Theta_{2}$. In special, when $\Theta_{1}=\Theta_{2}$ the system in Eq. 3 becomes undetermined. $f(\Theta_{1},\Theta_{2})$ reaches minimum at $\Theta_{1}=0$ and $\Theta_{2}=\pi/2$ or $\Theta_{1}=\pi/2$ and $\Theta_{2}=0$. However, as $\Theta_{1}$ or $\Theta_{2}$ approaches $\pi/2$, the length of the corresponding ion’s path at entry or exit increases, leading to greater systematic errors.
@@ 3 Results and discussion
To illustrate the application of the method, we calculated both the random and systematic uncertainties, along with the correlation factor, to optimize an experiment aiming at measuring the electronic stopping cross-section of gold for helium ions. The adopted film thickness in these calculations was 45 nm. The results for each contribution to the total uncertainty are presented in the panel of Fig. 2.
@@
The combination of random and systematic uncertainties yields the total uncertainties as shown in Fig. 3. We also calculated the final uncertainties for incident energies of 500 keV and 5000 keV, with the results presented in Figs. 4 and 5, respectively.
From these results, it is clear that, to keep the total uncertainty below 3% in the overall energy range, $\Theta_{1}$ should be kept between 50 and 80 and $\Theta_{2}$ should be below 30. Furthermore, it can be observed that the total uncertainties below 1% are not feasible in this experiment. If higher accuracy is needed, alternative methods should be explored [6].
@@
Fig. 6 shows the experimental data obtained using our method compared to SRIM and ICRU-49 data. The black points in the plot represent additional experimental data from the IAEA database. In the upper plot, we observe the overall agreement of our data with the semi-empirical models, while in the bottom plot, we display the relative residuals with SRIM as the reference. We provide densely spaced experimental data to better define the statistical limits of the method. At lower energies, it is evident that all experimental data fall within 4% in residuals amplitude, a value that is very close to the final uncertainty calculated using the method proposed here.
@@ 4 Conclusions
In this work, we have presented a novel methodology for optimization when measuring stopping power using backscattering data, offering an accurate and reliable approach for determining the stopping cross-section in thin films. Our method, which is based on a systematic treatment of uncertainties, provides a robust framework for minimizing random and systematic errors while optimizing geometries for precise measurements.

### v2

# Uncertainty minimization in electronic stopping cross-section measurements using the backscattering method

@@ Abstract
Accurate determination of electronic stopping cross sections is critical for ion beam analysis and related applications. While transmission methods are well established, backscattering approaches remain less explored from a metrological perspective, often lacking a systematic treatment of uncertainties. In this work, we present a quantitative framework to optimize experimental geometry in backscattering-based stopping measurements, explicitly accounting for both statistical and systematic errors. Applying the method to helium ions in gold thin films, we identify angular conditions that balance precision and accuracy, achieving total uncertainties below 3% over a wide energy range. The results, benchmarked against SRIM and ICRU-49, demonstrate that our approach improves the reliability of RBS-derived stopping data and strengthens their use for reference purposes and model validation.
@@
Stopping cross-section , Backscattering method , Uncertainty optimization , Systematic uncertainty
@@ 1 Introduction
The accurate determination of electronic stopping cross sections is essential for a broad range of applications in nuclear and materials science, including ion beam analysis (IBA), ion implantation, radiation damage studies, and detector development. The stopping power represents a fundamental quantity in the description of energy loss processes, and its precise knowledge is required both for modeling and for experimental quantification Jeynes [2017], Colaux et al. [2015]. Although theoretical understanding of the physical processes involved in energy loss phenomena is advancing Correa [2018], Zeb et al. [2013], Roth et al. [2017], De Vera et al. [2023], Matias et al. [2024a, b], experimental works focusing on traceable experimental measurements become crucial to benchmark models Moro et al. [2016]. Accurate experimental data also help to improve semiempirical approaches to produce tables that are used in simulations targeting a wide range of applications.
Historically, several experimental approaches have been developed to measure stopping cross sections, the most established being the transmission method and the Rutherford backscattering spectrometry (RBS)-based method Bauer [1987], Mertens [1987]. While the transmission method is often regarded as more direct, it is not exempt from systematic uncertainties, including foil non-uniformity, energy calibration issues, and beam stability Moro et al. [2016], Mertens and Semrad [1986]. Moreover, the perception that it offers intrinsically higher accuracy has contributed to an imbalance in the literature regarding the metrological scrutiny of alternative methods.
In particular, despite its wide use in IBA, the backscattering method remains relatively underexplored from a metrological standpoint. Although early works by Bauer Bauer [1987] and Mertens Mertens and Semrad [1986] have laid important foundations for the method, and some efforts have been made to evaluate the accuracy of RBS-derived stopping data Bergsmann and Bauer [2001], Paul and Schinner [2004], Pascual-Izarra et al. [2005], recent literature still lacks systematic frameworks to quantify and minimize both statistical and systematic uncertainties in these measurements. As a consequence, experimental stopping cross-section data obtained via RBS often omit a complete uncertainty budget, which compromises their traceability and utility as reference data Colaux et al. [2015].
There are different methodologies and strategies to determine experimentally the stopping cross-section in backscattering experiments. Some are based on relative measurements to tables of other better-known elements Andersen et al. [1978], Meersschaut et al. [2024], some are constrained by a fitting formula Eppacher and Semrad [1988], Pascual-Izarra et al. [2005], and others need some thickness markers Leblanc and Ross [1994].
This work aims to address this gap by providing a detailed optimization strategy for stopping cross section measurements using the backscattering method. We present a methodology that identifies optimal geometrical configurations to minimize total uncertainty, incorporating both random and systematic components. Our approach builds upon earlier formulations Bauer [1987], Bergsmann and Bauer [2001], but extends them with explicit error propagation, correlation analysis, and benchmarking against reference models such as SRIM and ICRU tables. The methodology is applied to helium ions backscattered from thin gold films, for which precise reference data are available.
By offering a reproducible and uncertainty-aware approach to stopping power determination via backscattering, we contribute to improving the reliability of this technique and enabling its broader adoption in precision applications.
@@ 2.1 The back-scattering method
There are several approaches to measuring stopping cross-section using backscattering data Bauer [1987]. In this work, we adopt the method based on determining the peak width in thin-film measurements, as described in Chu et al. [1978]. Fig. 1 shows a typical spectrum of a pure metallic film deposited on a light substrate. The peak is formed by all scattered ions from the surface down to the film-substrate interface. Its width ($\Delta E$) is the difference between the energy of ions scattered by the film atoms at the surface ($K\cdot E_{0}$, with $K$ being the kinematic factor and $E_{0}$ the incidence beam energy) and the exit energy of ions re-emerging after being scattered by film atoms at the interface ($E_{out}$). Assuming a homogeneous stopping in the test film along both, the entry and exit paths, the energy difference can be written as:
@@
The peak width can be determined by adding the energy loss during the path of the ion into the film and the energy loss during its exit from a thin film of thickness $t$. Assuming homogeneous stopping along both the entry and exit paths, this expression can be written as:
@@
being the stopping cross section in the $j$ path, either the entry (index $in$) or the exit (index outout) of the film. $\theta_{in}$ and $\theta_{out}$ represent the angles between the ion’s entry and exit paths and the surface normal, respectively. $N$ is the material atomic density.
The fundamental principle of this methodology is that by measuring the peak width in two different geometrical configurations (i.e. two pairs of values of $\theta_{in}$ and $\theta_{out}$) for the same incident energy, one can obtain a system of equations. The solutions for $\epsilon_{in}]$ and $\epsilon_{out}]$ yield the stopping cross section at two different mean energies. In matrix formalism, this can be expressed as:
@@
where index $A$ and $B$ denote the two different measurements. To better constrain the problem, we impose the conditions $\theta_{in}^{A}=\theta_{out}^{B}=\Theta_{1}$ and $\theta_{in}^{B}=\theta_{out}^{A}=\Theta_{2}$. This assumption is necessary to reduce the number of variables, thereby enhancing result visualization and simplifying the experimental procedure.
Two things are important to note: the first is that if $\Theta_{1}=\Theta_{2}$ the equation system of Eq. 3 becomes undetermined; and the second is that the factor $K$ implicitly depends on the scattering angle, which remains fixed throughout the analysis. As a consequence, the experimental procedure involves tilting the sample in such a way that, in the first measurement, the incidence angle is $\Theta_{1}$ and the exit angle is $\Theta_{2}$, while in the second measurement, the angles are reversed, with the incidence angle set to $\Theta_{2}$ and the exit angle to $\Theta_{1}$.
There are different approximations to determine the mean energy at the entry and exit paths of the film Chu et al. [1978]. Here, we adopt the following approximations:
@@
with $\Delta\bar{E}=(\Delta E^{A}+\Delta E^{B})/2$. This is an alternative definition based on the symmetrical mean energy approximation proposed in Chu et al. [1978], which is more convenient for an experimental approach.
@@ 2.2 Accounting for random uncertainties
The random uncertainty in this work arises primarily from the counting statistics of the backscattering spectra, which determine the precision with which the edges of the thin-film peak can be located. In practice, this translates into an uncertainty associated with the discretization of the spectrum into channels, where a uniform probability distribution is assumed within each channel when evaluating the peak width ($\Delta E$).
From Eq. 3, we can propagate the uncertainties in the determination of $\Delta E_{j}$ to the values of $\epsilon_{j}]$ using the definition of a covariance matrix Helene et al. [2016]:
@@
$X=\begin{pmatrix}\frac{K\cdot t}{\cos(\Theta_{1})}&\frac{t}{\cos(\Theta_{2})}\\ \\ \frac{K\cdot t}{\cos(\Theta_{2})}&\frac{t}{\cos(\Theta_{1})}\\ \end{pmatrix}\ \ \ \ \$(6)
@@
and the outer diagonal results in the covariance between the $\epsilon]_{in}$ and $\epsilon]_{out}$ values, which is used to calculate the correlation factor r as:
@@ 2.3 Accounting for systematic error
In this work, the systematic error refers to the deviation introduced by approximations used in the analytical method, such as assuming a constant stopping cross section along the ion path and estimating the mean energy using simplified expressions. These deviations are quantified by comparing the analytical results with reference values obtained from numerical integration of the SRIM stopping cross-section data.
To estimate the systematic uncertainties, we tested Eqs. 4 and 8 against the SRIM curve. The peak width, $\Delta E$, has been computed directly using the integral (as in Eq. 12) form and SRIM data, while the methodology described in Section 2.1 is applied to determine $\epsilon]$ using Eq. 8. These values are then compared to the SRIM values themselves interpolated at the mean energies given by Eq. 4. The differences between the interpolated and the output of the matrix calculation are considered as systematic errors. This accounts for systematic errors inserted either by the surface approximation or by the calculation of the mean energy in the entry and exit paths.
$\begin{split}E_{1}&=E_{0}-\int_{0}^{t/\cos(\theta_{in})}{\left[\frac{1}{N}\frac{dE}{dx}(E(x),x)\right]_{E(0)=E_{0}}dx}\\ E_{out}&=K\cdot E_{1}-\int_{0}^{t/\cos(\theta_{out})}{\left[\frac{1}{N}\frac{dE}{dx}(E(x),x)\right]_{E(0)=K\cdot E_{1}}dx}\\ \Delta E&=K\cdot E_{0}-E_{out}\end{split}$ (12)
@@
with $s({[\epsilon]_{j}})$ being the systematic error and $\textbf{SRIM}(\bar{E_{j}})$ being the SRIM interpolation at the energy $\bar{E_{j}}$. $j$ is an index that denotes $in$ or outout.
2.4 Other uncertainty sources
In addition to the statistical and systematic errors explicitly addressed in this work, several other factors contribute to the overall uncertainty in stopping cross-section measurements. These include uncertainties related to the angular alignment of the beam and detector, variations in the material density, energy calibration, detector resolution, and the intrinsic energy straggling of the ions.
Energy-loss straggling also contributes to the broadening of the backscattering peak. Although not explicitly included in our analytical propagation, its effect largely follows the same trend as the systematic deviations discussed in Section 2.3, i.e., it increases with the effective path length of ions inside the film. Since the aim of the present work is not to provide absolute uncertainty values but rather a metric to optimize measurement geometries, minimizing the systematic error in practice also minimizes the impact of straggling. In this sense, the optimization strategy remains valid even without an explicit quantitative treatment of straggling.
While a comprehensive treatment of all these effects is beyond the scope of this article, we focus here on developing a strategy for optimizing the geometric configuration of the measurement setup. By propagating selected, dominant sources of uncertainty through our analytical framework, namely, the uncertainty in the energy width ($\Delta E$), we identify angular combinations that minimize the propagated error. A key feature of this approach is the explicit treatment of the correlation between entry and exit path measurements, a factor that is often neglected in conventional stopping cross-section evaluations, yet plays a critical role in minimizing combined uncertainties.
An additional and important source of uncertainty arises from the characterization of the film thickness. While this quantity must be accurately determined using independent experimental techniques (e.g., profilometry or Rutherford Backscattering Spectrometry), its uncertainty should be quadratically summed to the final error obtained by our method. It is important to highlight, however, that the absolute value of the film thickness not only contributes as a multiplicative factor to the stopping cross section determination, but also directly affects the propagation of both statistical and systematic errors, as discussed in Section 2.2 and 2.3. This effect is particularly significant for the systematic component, due to its dependence on the energy loss along the ion trajectory.
@@ 2.5 Finding the best geometry
With the definitions of random and systematic uncertainties for the case of determination of $\epsilon]$, the total uncertainty can be defined as:
@@
Note that the common factor in the statistical uncertainty of $\epsilon]_{in}$ and $\epsilon]_{out}$, as in Eq. 9 and 10, can be written as the function $f(\Theta_{1},\Theta_{2})$:
@@
which diverges when $\Theta_{1}\approx\Theta_{2}$. In special, when $\Theta_{1}=\Theta_{2}$ the system in Eq. 3 becomes undetermined. $f(\Theta_{1},\Theta_{2})$ reaches a minimum at $\Theta_{1}=0$ and $\Theta_{2}=\pi/2$ or $\Theta_{1}=\pi/2$ and $\Theta_{2}=0$. However, as $\Theta_{1}$ or $\Theta_{2}$ approaches $\pi/2$, the length of the corresponding ion’s path at entry or exit increases, leading to greater systematic errors.
@@ 3 Results and discussion
To illustrate the application of the method, we calculated both the random and systematic uncertainties, along with the correlation factor, to optimize an experiment aiming at measuring the electronic stopping cross-section of gold for helium ions. The adopted film thickness in these calculations was 45 nm ($\approx 266\times 10^{15}$ at./cm2).
Observing Eqs. 3 and 12, both the random and systematic uncertainties depend on the film thickness $t$. In this context, the validity of the method relies on representative measurements of $\Delta E$. This imposes an upper limit on the incident beam energy, as it becomes constrained by the achievable depth resolution—i.e., when $\Delta E$ approaches the detector resolution, the depth information is no longer reliable. Assuming a detector resolution of 12 keV (FWHM), depth resolution calculations using ResolNRA Mayer [2008] indicate that a 45 nm film is above this limitation within our energy range of 500-5100 keV.
Figure 2: Depth-resolution calculated using ResolNRA demonstrating the 45 nm thick film is suitable for stopping cross-section measurements in the energy range of this study.
The results for each contribution to the total uncertainty are presented in the panel of Fig. 3.
@@
The combination of random and systematic uncertainties yields the total uncertainties as shown in Fig. 4. We also calculated the final uncertainties for incident energies of 500 keV and 5000 keV, with the results presented in Figs. 5 and 6, respectively.
From these results, it is clear that, to keep the total uncertainty below 3% in the overall energy range, $\Theta_{1}$ should be kept between 50 and 80 and $\Theta_{2}$ should be below 30. Furthermore, it can be observed that the total uncertainties below 1% are not feasible in this experiment. If higher accuracy is needed, alternative methods should be explored Moro et al. [2016].
@@
Fig. 7 shows the experimental data obtained using the method described in 2.1 with a combination of angles chosen accordingly to our approach. The results are compared to SRIM and ICRU-49 data. The black points in the plot represent additional experimental data from the IAEA database. In the upper plot, we observe the overall agreement of our data with the semi-empirical models, while in the bottom plot, we display the relative residuals with SRIM as the reference. We provide densely spaced experimental data to better define the statistical limits of the method. At lower energies, it is evident that all experimental data fall within 4% in residuals amplitude, a value that is very close to the final uncertainty calculated using the method proposed here.
@@ 4 Conclusions
In this work, we have presented a novel methodology for optimization when measuring stopping cross-section using backscattering data, offering an accurate and reliable approach for determining the stopping cross-section in thin films. Our method, which is based on a systematic treatment of uncertainties, provides a robust framework for minimizing random and systematic errors while optimizing geometries for precise measurements.
It is worth noting that energy straggling, although not explicitly included, is expected to follow the same angular dependence as the systematic uncertainties, and thus its effect is indirectly minimized by the present optimization strategy.